# SelectLLM: Query-Aware Efficient Selection Algorithm for Large Language Models

### v1

@@ Appendix A Prompting Templates and Answer Extraction
Considering diverse LLMs and benchmarks adds challenges to prompting as no single uniform prompting approach best fits all LLMs (Sclar et al., 2023). Based on the insights from recent work on the appropriate usage of prompts (Sahoo et al., 2024) and from our own experiments, we make the following observations about the prompting trends: (1) For non-chat LLMs, few-shot chain-of-thought (COT) (Wei et al., 2022) prompting works better than zero-shot (Kojima et al., 2022) across both datasets, so we use 5 few-shot random examples obtained from the development set. The few-shot prompting leads to over 95% viable answers (except for llama2-7b-lm LLM which has the viability score of 83%) in generated solutions, where we consider an answer to be viable if it is represented by a single numeric/alphabetic string that can be extracted from the generated solution to compare with the reference answer. At the same time, a viable answer can be correct or incorrect. Viability is estimated using an automated script and is verified by manual inspection. (2) For chat LLMs, few-shot COT distracts the generator, which leads to unexpected outputs, so the zero-shot COT works best. To ensure correctness, we utilize different models’ chat templates from HuggingFace. [^1] https://huggingface.co/docs/transformers/en/chat_templating The viability of answer extraction for chat models is $\sim$92%. Examples of zero-shot and few-shot prompting are presented in Appendix Figure 4.
@@
Table 4: Statistics on the datasets: For the MMLU dataset, we have swapped the officially released training and test splits to achieve a similar distribution as the GSM8K dataset.
LLMs Chat? Spec? #params
@@
Table 6: Runtime statistics on various LLMs over 10 generations for each input query. The timings are recorded using a single A100 GPU. ‘sec’ denotes seconds, and COT denotes Chain-of-thought. For few-shot COT, we have considered 5 random examples from the development set.
Figure 4: Sample GSM8K zero-shot COT prompt for a chat (or instruction-tuned) LLMs, and MMLU few-shot COT prompt for non-chat LLMs.
The adapted prompting approaches used in our LLM queries are designed to instruct LLMs to specify that their final answers should be provided at the very end of each of their responses. We thus use a simple answer extraction policy of selecting the last mentioned numerical value (for GSM8K) and multiple-choice option (for MMLU) from the generated responses. Figure 4 in the Appendix shows a sample generation example. Responses failing to include any final answer are considered invalid (‘INVALID’) and counted as incorrect responses. For MMLU, we evaluate the extracted options directly against the annotated correct answers (‘A’, ‘B’, ‘C’, and ‘D’) from the dataset. For GSM8K, questions where the absolute difference between the ground truth and predicted numerical answers is less than $\epsilon=0.1$ are evaluated as solved correctly. This threshold was set to accommodate instances where model-generated real-valued answers differ slightly from the expected answers, e.g., due to rounding errors.
@@ Appendix B Data Statistics and Selected LLMs
The statistics of the dataset and the list of considered diverse LLMs are presented in Table 5 and 5, respectively.
@@ Appendix D Sample Prompts
Appendix E Distribution of Different LLMs in Top-$s$ Subset Selected with SelectLLM
@@
Appendix F Query Response Distribution with Different Models
@@
Appendix G Implementation Details
@@
We use the vLLM [^2] https://github.com/vllm-project/vllm package to query LLMs. All models were queried with a temperature of $0.8$ and a max token length of $2000$. Each question prompt was queried 10 times with different initialization seeds. We used a single NVIDIA A100 GPU for all runs. Querying each dataset once took approximately 1-2 hours.
@@
We use the HuggingFace [^3] https://huggingface.co/ library for loading and tuning all pre-trained Transformer encoders in our experiments. Each model was trained for $10$ epochs, with an initial learning rate of $1e$-$6$, a warmup ratio of $0.1$, and class-balanced CrossEntropy loss. The training checkpoint with the lowest validation loss was selected for inference.
Appendix H Ablation Studies
@@
Appendix I Importance of Domain: A Case Study with MMLU

### v3

@@ Appendix A Prompting Templates and Answer Extraction
Considering diverse LLMs and benchmarks adds challenges to prompting, as no single uniform prompting approach fits all LLMs (Sclar et al., 2023). Based on insights from recent work on the appropriate usage of prompts (Sahoo et al., 2024) and our own experiments, we make the following observations about prompting trends:
  1. 1.
    For non-chat LLMs, few-shot chain-of-thought (COT) prompting (Wei et al., 2022) works better than zero-shot (Kojima et al., 2022) across both datasets. Therefore, we use five few-shot random examples obtained from the development set. The few-shot prompting results in over 95% viable answers (except for the llama2-7b-lm LLM, which has a viability score of 83%). An answer is considered viable if it is represented by a single numeric/alphabetic string that can be extracted from the generated solution to compare with the reference answer. Viability is estimated using an automated regular expressions based-script and verified through manual inspection.
  2. 2.
    For chat LLMs, few-shot COT distracts the generator, leading to unexpected outputs, so zero-shot COT works best. To ensure correctness, we utilize different models’ chat templates from HuggingFace. [^2] https://huggingface.co/docs/transformers/en/chat_templating The viability of answer extraction for chat models is approximately 92%. Examples of zero-shot and few-shot prompting are presented in Appendix Figure 4.
The limitation of answer extraction from 92% to 95% (except for llama2-7b-lm) can be attributed to two factors: (i) limitations in the LLMs’ response generation, where outputs are not structured in an extractable format, and (ii) limitations in the extraction algorithm, which fails to accurately parse the generated text. For more details, see the limitations section.
@@
Table 5: Dataset statistics for GSM8K and MMLU benchmarks. For the MMLU, the officially released training and test splits have been swapped to align with the distribution of the GSM8K.
LLMs Chat? Spec? #Parameters
@@
Table 7: Runtime statistics on various LLMs over M generations for each input query. The timings are recorded using a single A100 GPU. ‘sec’ denotes seconds, and COT denotes Chain-of-thought. For few-shot COT, we have considered 5 random examples from the development set.
Figure 4: Sample zero-shot COT prompt for a chat (or instruction-tuned) LLM is shown for a GSM8K sample, and the few-shot COT prompt for non-chat LLMs is shown for an MMLU sample. Note that the type of prompting is associated with the LLM type rather than the dataset. Specifically, non-chat LLMs use few-shot COT, while chat models use zero-shot COT.
The adapted prompting approaches used in our LLM queries are designed to instruct LLMs to specify that their final answers should be provided at the very end of each of their responses. We thus use a simple answer extraction policy of selecting the last mentioned numerical value (for GSM8K) and multiple-choice option (for MMLU) from the generated responses. Figure 4 in the Appendix shows a sample generation example. Responses failing to include any final answer are considered non-viable or invalid (‘INVALID’) and counted as incorrect responses. For MMLU, we evaluate the extracted options directly against the annotated correct answers (‘A’, ‘B’, ‘C’, and ‘D’) from the dataset. For GSM8K, questions where the absolute difference between the ground truth and predicted numerical answers is less than $\epsilon=0.1$ are evaluated as solved correctly. This threshold was set to accommodate instances where model-generated real-valued answers differ slightly from the expected answers, e.g., due to rounding errors.
@@ Appendix B Data Statistics and Selected LLMs
The statistics of the dataset and the list of considered diverse LLMs are presented in Table 6 and 6, respectively.
@@ Appendix D Sample Prompts
Appendix E SelectLLM Inference Algorithm
Refer to the Algorithm 1.
[…]
3:for each query $q_{i}$ in $\mathcal{Q}$ do
7:     Initialize an empty set for answer set $A_{i}\leftarrow\phi$
8:     for each $l_{j}$ in $L_{s}$ do
9:         Generate M responses: $a_{i}^{M}=l_{j}(q_{i})$triangleright Generate M responses with LLM $l_{j}$
10:         Find answer frequency $a_{i}^{f}=\{a_{i}^{k}:\text{countof}(a_{i}^{k})/\sqrt{c_{i}^{(s)j}}\,|\,\text{for }a_{i}^{k}\text{ in unique}(a_{i}^{K})\}$
11:         $A_{i}\leftarrow A_{i}\cup a_{i}^{f}$
12:     end for
13:     Return most frequent answer in $A_{i}$
14:end for
Model/Setup GSM8K MMLU
BERT (base-uncased) 0.63 0.59
BERT (base-uncased) + Weighted-label penalty 0.64 0.62
RoBERTa (base) 0.67 0.61
BRoBERTa (base) + Weighted-label penalty 0.71 0.68
T5 (base) 0.66 0.60
T5 (base) + Weighted-label penalty 0.69 0.65
Table 8: The F1 score for the MLC classifier is evaluated with different pre-trained language models. Additionally, a weighted-label penalty is applied to handle label imbalance, inspired by Zhang et al. (2020). This evaluation is conducted across both GSM8K and MMLU datasets.
Appendix F Distribution of Different LLMs in Top-s Subset Selected with SelectLLM
@@
Appendix G Query Response Distribution with Different Models
@@
Appendix H Implementation Details
@@
We use the vLLM [^3] https://github.com/vllm-project/vllm package to query LLMs. All models were queried with a temperature of $0.8$ and a max token length of $2000$. Each question prompt was queried K times with different initialization seeds. We used a single NVIDIA A100 GPU for all runs. Querying each dataset once took approximately 1-2 hours.
@@
We use the HuggingFace [^4] https://huggingface.co/ library for loading and tuning all pre-trained Transformer encoders in our experiments. Each model was trained for $10$ epochs, with an initial learning rate of $1e$-$6$, a warmup ratio of $0.1$, and class-balanced CrossEntropy loss. The training checkpoint with the lowest validation loss was selected for inference. In this study, we have considered M=10 inspired by Li et al. (2024).
Figure 7: Latency vs. accuracy vs. $s$-values plots for different SelectLLM policies on GSM8K (left) and MMLU (right) validation sets. $s$-value: total number of LLMs selected with SelectLLM.
Appendix I Latency vs. Accuracy vs. s-Value:
Figure 7 shows the relationship between latencies, accuracies, and $s$-values for different SelectLLM policies across both datasets with corresponding development set. WeightedMaxConf consistently outperforms MaxConf and WeightedMaxConf on both accuracy and latency for most $s$-values (except 5 and 6 for GSM8K). This indicates the superiority of WeightedMaxConf, maintaining effectiveness even with a small number of LLMs. A larger number of LLMs and higher latency are required for GSM8K, while a lower $s$-value and latency are needed for MMLU to achieve high accuracy. Selecting 4 LLMs for GSM8K and 2 for MMLU achieves optimal performance, indicating the effectiveness and efficiency of SelectLLM with WeightedMaxConf. The results are presented on the validation set, and the optimal value of $s$ is obtained. Using this optimal value, the results are reported on the test set in Table 2.
Appendix J Performance of Individual LLMs
Table 9 presents the individual LLM performance for both the GSM8K and MMLU datasets. There is no clear best-performing model for both datasets. For instance, the gemma-7b-lm model performs the best in terms of accuracy for GSM8K, but its latency cost is also high. Similar trends are observed for the MMLU dataset. The performance of individual LLMs should not be directly compared with the SelectLLM model. Selecting the best LLM with high accuracy from a large pool requires running inferences across all LLMs to identify the top performer, resulting in the cumulative latency of all models, rather than just the selected one. On the other hand, randomly selecting an LLM may not necessarily choose the best performer. In this context, comparing SelectLLM with ensemble-based baselines is more appropriate, which is done in the main results (Table 2) where SelectLLM outperforms all baseline models, such as strange LLM-Blender baseline, in terms of both latency and accuracy.
Models / Setups GSM8K MMLU
Acc $\uparrow$ Lat $\downarrow$ Acc $\uparrow$ Lat $\downarrow$
gemma-7b-lm 71.27 7.10 63.73 3.00
mistral-7b-lm 60.50 3.70 61.57 1.80
metamath-7b-lm 67.25 4.70 41.76 2.40
llama2-7b-lm 48.10 2.30
llama2-13b-chat 49.20 1.80 52.94 4.80
mistral-7b-it 56.71 1.00 53.92 1.10
gemma-7b-it 42.23 0.70 50.72 1.00
Table 9: Performance and latency scores for different LLMs on GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021) test sets. Acc: accuracy in percentage (%); Lat: latency in seconds; ‘–’: values are not available.
Appendix K Ablation Studies
@@
Appendix L Importance of Domain: A Case Study with MMLU
@@
Appendix M SHAP Analysis
In this section, we provide the details of the SHAP-values-based qualitative analysis discussed in Section 6.1.
Extracting Shapley Values
We first extract token-level Shapley values for the input question text to the fine-tuned MLC classifiers for each LLM using the PartitionExplainer module of the SHAP library  [^5] https://shap.readthedocs.io/en/latest/. The token-level SHAP values are then summed according to word boundaries to get word-level SHAP values. This is a valid operation as SHAP values are additive in nature.
Synset Mapping
In order to interpret the impact of different semantic word categories on the predicted solvability of questions, we map all viable words to their corresponding synsets using NLTK’s WordNet module  [^6] https://www.nltk.org/. In this process, all instances of a word are recorded as instances of its respective synset. Next, we merge entries for synsets that represent less than 0.3% (decided by observation) of the total number of tokens with their hypernyms. This helps reduce sparsity in the synsets to review. The mean SHAP values for each synset are then calculated across the test set of the datasets.
Synset Grouping
Based on the top 15 most detrimental synsets for each LLM in each of the four sets (i.e., having the lowest non-positive average SHAP values), we manually identify groups of related and recurring synsets. The groups for GSM8K, as presented in Table 3, are as follows:
  • frequency terms:
    Synset(’twice.r.01’), Synset(’thrice.r.01’), Synset(’times.n.01’), etc.
  • time duration units:
    Synset(’hours.n.01’), Synset(’day.n.01’), Synset(’week.n.01’), Synset(’time_unit.n.01’), Synset(’time_period.n.01’), etc.
  • age units: Though similar to time-units, these synsets correspond specifically to instances where the age of an entity is being described. These include: Synset(’age.n.01’), Synset(’year.n.01’), Synset(’time_of_life.n.01’), etc.
  • small numbers: Though some small integers and numbers posses their own individual synsets, we identify and filter all instances of small numbers ($\leq$100) (in numerical and word form) before synset matching and group them as small numbers.
  • quantifiers:
    Synset(’every.s.01’), Synset(’many.a.01’), Synset(’more.a.01’), Synset(’less.a.01’), etc.
  • ordinals:
    Synset(’first.a.01’), Synset(’second.s.01’), Synset(’third.s.01’), etc.
  • fractional values:
    Synset(’common_fraction.n.01’).
  • rates and ratios:
    Synset(’rate.n.01’), Synset(’ratio.n.01’), Synset(’proportion.n.01’), Synset(’magnitude_relation.n.01’), etc.
  • named entities: As WordNet does not record proper-nouns or named-entities like persons, organizations, locations, dates, and times, we identify such words before synset matching and designate them a separate group.
  • other units:
    Synset(’monetary_unit.n.01’), Synset(’linear_unit.n.01’), Synset(’work_unit.n.01’), Synset(’area_unit.n.01’), Synset(’definite_quantity.n.01’), etc.
  • groups:
    Synset(’family.n.01’), Synset(’team.n.01’), Synset(’unit.n.03’), etc.